Q\Determine the following ODEs are exact or not? Why?
((y)dx+(x)dy/(x+y)^2)+dy=0
1
Expert's answer
2020-10-28T18:47:40-0400

"\displaystyle

y\mathrm{d}x + \frac{x\mathrm{d}y}{(x+y)^2} + \mathrm{d}y = 0\\


\textsf{Multiplying through by}\, (x + y)^2\\

y(x + y)^2\mathrm{d}x + x\mathrm{d}y + (x + y)^2\mathrm{d}y = 0\\


y(x + y)^2\mathrm{d}x + (x + (x + y)^2)\mathrm{d}y = 0\\

M = y(x + y)^2\\

\begin{aligned}
M_y = \frac{\partial M}{\partial y} &= (x + y)^2 + 2y(x + y)\\
&= (x + y)(x + y + 2y)\\
&= (x + 3y)(x + y)
\end{aligned}\\

N = x + (x + y)^2\\

N_x = \frac{\partial N}{\partial x} = 1 + 2(x + y)\\

\textsf{Since}\, N_x \neq M_y,\, \textsf{the differential equation}\\
\textsf{is not exact.}"